OpenStudy (anonymous):

Lee needed to make drapes the length from the floor to the top of the window. The wall beneath the window had the height of 2 2/3 feet and the height of just the window measured 2 3/4 feet. How long do the drapes have to be?

OpenStudy (anonymous):

2 2/3 + 2 3/4= 5 5/12 2 2/3=8/3=32/12 2 3/4=11/4=33/12 32/12+33/12=65/12=5 5/12 ft